OpenStudy (anonymous):

Can someone please help me with this question? Find Pk+1 for the given Pk.

OpenStudy (anonymous):

\[P _{k}=k^2(2k+2)\]

OpenStudy (anonymous):

Replace every \(k\) you see with \(k+1\): \[P_k:~k^2(2k+2)~~\Rightarrow~~P_{k+2}:~(k+1)^2(2(k+1)+2)\]

OpenStudy (anonymous):

Then what? I dont understand, I am sorry.

OpenStudy (anonymous):

Well, that would be an acceptable answer, though you could simplify it a bit. \[(k+1)^2(2(k+1)+2)=(k+1)^2(2k+4)\]

OpenStudy (anonymous):

Alright, thank you so much. I appreciate your help!